Circular No. 8352

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 L1 (LINEAR) An apparently asteroidal object discovered by the LINEAR project

(discovery observation below), and posted on the ‘NEO Confirmation Page’, has been found to show cometary appearance (5′′ round coma without condensation) by J. Young (0.6-m reflector, Table Mountain) on CCD images taken on June 14.3–14.4 UT.

2004 UT  α_{2000}  δ_{2000}  Mag.
June 12.27759  17^{h}58^{m}16.^{s}96  −12^{o}54′06′′.6  18.0

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2004-L58.

T = 2005 Mar. 31.408 TT  ω = 239.^{o}431
Ω = 65.634 }2000.0
q = 2.18138 AU  i = 159.771

SUPERNOVAE 2004cd, 2004ce, 2004cf, 2004cg M. Riello, European Southern Observatory (ESO), reports (together

with S. Benetti, E. Cappellaro, F. Patat, M. T. Botticella, M. Turatto, and S. Valenti) the discovery of four supernovae on CCD frames taken on May 14 during a search performed with the Max-Planck-Institut 2.2-m telescope (+ Wide Field Imager) at ESO.

SN  2004 UT  α_{2000}  δ_{2000}  R  z Type
2004cd May 14 13^{h}44^{m}50.^{s}70 − 0^{o}03′48′′.3 22.7 0.5 Ia
2004ce May 14 13 54 45.93 −12 14 36.3 22.8 0.465 Ia
2004cf May 14 14 11 05.77 −11 44 09.4 22.9 0.247 Ib/c
2004cg May 14 14 11 23.71 −12 01 08.4 22.4 0.264 II

Spectroscopic classification was based on observations (range 600–1100 nm) obtained on May 24 with the ESO 8.2-m YEPUN telescope (+ FORS2). Nothing was visible at the above positions on reference images taken with the same equipment in 2001 and 2002 (limiting mag R ∼ 23.5). SNe 2004ad, 2004ae, and 2004ag are all < 0′′.2 from their host-galaxy centers; SN 2004cf is 1′′.0 west and 1′′.0 south of its host-galaxy center.

COMET C/2004 K3 (LINEAR) Corrigendum. On IAUC 8350, line 6, for Holvercem read Holvorcem

2004 June 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green